OpenStudy (nuccioreggie):

Line Q is represented by the following equation: 2x + y = 11 Which equation completes the system that is satisfied by the solution (3, 5)? x + 2y = 15 x + y = 15 x + y = 8 x − y = 2

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (whovianchick):

Aight this is my last one though I've actually got school to finish today XD

OpenStudy (whovianchick):

This one is pretty simple. You know that the first equation has a solution of (3,5), you just need to figure if another one does, too. We do that by plugging it in. x + 2y = 15 (3) + 2(5) = 15 NO 3+10=15 13=15 x + y = 15 (3) + (5) = 15 NO 8= 15 x + y = 8 (3) + (5) = 8 YES 8=8 x − y = 2 (3)-(5)=2 NO -2=2

OpenStudy (whovianchick):

The answer is C

OpenStudy (whovianchick):

And if you only have 3 left I'll see what I can do

OpenStudy (nuccioreggie):

ok (: